# SPHERE: Semantic-PHysical Engaged REpresentation for 3D Semantic Scene Completion

### v1

@@ 1. Introduction
In autonomous driving systems, 3D Semantic Scene Completion (SSC)  (Roldao et al., 2022) is a crucial task for holistic scene perception, inferring voxel-level geometry and semantics for several downstream tasks. To deal with the inherent complexity of large-scale outdoor scenarios, LiDAR-based SSC methods (Roldao et al., 2020; Cheng et al., 2021; Rist et al., 2021; Yan et al., 2021; Xia et al., 2023; Zhang et al., 2023a; Ye et al., 2022; Li et al., 2023e; Lee et al., 2024) have been the predominant solutions, taking advantage of explicit 3D geometric structural information entailed in the point cloud data. Despite their effectiveness, the inherently high cost and limited scalability of LiDAR sensors constrain practical applications of LiDAR-based methods in real-world deployment. In light of this issue, multi-view images (Miao et al., 2023; Yao et al., 2023; Tong et al., 2023; Zheng et al., 2024) have garnered increasing attention as a more cost-effective and flexible input modality.
The pioneering camera-based SSC method, MonoScene (Cao and De Charette, 2022), extracts 2D features from input RGB images and lifts them into dense 3D volume features through depth-based projection. Following this, existing camera-based SSC methods fall mainly into two paradigms: voxel-based and plane-based methods. Voxel-based SSC methods (Zhang et al., 2023b; Li et al., 2023d) discretize 3D scenes into voxels and assign extracted image features to each voxel, and progress has been achieved in exploiting scene-from-instance feature interactions (Jiang et al., 2024), and employing a dense-sparse-dense strategy for hybrid guidance (Mei et al., 2024). In addition, plane-based SSC methods adopt Bird’s-Eye-View (BEV) (Huang et al., 2021) and Tri-Perspective View (TPV) (Huang et al., 2023) features for efficient scene representations. Subsequent research further improves the quality of scene representation through camera-aware depth estimation and refinement (Li et al., 2023b), dynamic temporal stereo information (Li et al., 2023a), and explicit-implicit compact feature projection (Ma et al., 2024).
While existing SSC methods have demonstrated promising performance in voxel-wise semantic prediction, they frequently fail to capture the underlying physical regularities in the real world, leading to unrealistic geometric details in scene completion results, as illustrated in Figure 1 (a). On the other hand, neural reconstruction methods like NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023) excel at learning realistic geometry details, but suffer from high computational expense and slow convergence in large-scale complex autonomous driving scenes, resulting in confusing semantics, as shown in Figure 1 (b). The above challenges naturally raise a critical question: How can we achieve both accurate semantics and realistic geometry, efficiently and effectively?
@@ 2.1. Camera-based Semantic Scene Completion
Voxel-based SSC methods discretize 3D space into voxels and assign image features to each voxel for scene perception. MonoScene (Cao and De Charette, 2022) is a pioneering effort that reconstructs 3D scenes from RGB images by projecting image features to 3D space along optical rays. OccFormer (Zhang et al., 2023b) introduces a dual-path transformer architecture with a mask-wise prediction paradigm, effectively capturing spatial dependencies for improved completion accuracy. VoxFormer (Li et al., 2023d) presents a two-stage framework that integrates class-agnostic query proposals with class-specific semantic segmentation, facilitating the propagation of discriminative semantic information from seed voxels across the entire scene for enhanced completion. Symphonize (Jiang et al., 2024) introduces a scene-from-instance framework that enhances the interplay between image and volumetric representations, enabling more precise scene completion. SGN (Mei et al., 2024) presents a one-stage SSC framework employing a dense-sparse-dense design, which enhances segmentation boundary clarity and improves the accuracy of 3D semantic scene completion. CGFormer (Yu et al., 2024) introduces a context- and geometry-aware voxel transformer that utilizes context-dependent queries and multiple 3D representations to enhance scene completion performance.
Plane-based SSC methods compress the 3D space along certain dimensions for efficient view transformation and scene representation. A pioneering approach in this domain is the Lift-Splat-Shoot (LSS) (Philion and Fidler, 2020) framework, which establishes an end-to-end pipeline by first ”lifting” images into feature frustums, then ”splatting” these frustums onto a rasterized BEV grid, and finally ”shooting” template trajectories into a BEV cost map for downstream tasks. Building upon the LSS paradigm, BEVDet (Huang et al., 2021) introduces a holistic BEV-based framework for scene understanding, comprising an Image-view Encoder, a View Transformer, a BEV Encoder, and a Task-specific Head. Subsequent achievements have been made in further enhancing the BEV-based representations, incorporating techniques such as inverse perspective mapping (Hu et al., 2021), angle and radius aware rasterization (Liu et al., 2023), and cross-attention layers with BEV queries (Li et al., 2022). BEVDepth (Li et al., 2023b) enhances depth learning by introducing a camera-aware depth estimation module alongside a depth refinement module, improving the accuracy of depth predictions. Building on this, BEVStereo (Li et al., 2023a) advances depth estimation by incorporating dynamic temporal stereo information. Beyond BEV, TPVFormer (Huang et al., 2023) introduces a Tri-Perspective View (TPV) representation, which decomposes voxel grids into three orthogonal planes, allowing efficient scene encoding.
@@ 2.2. Neural Representation in 3D Perception
Neural representation techniques have emerged as powerful tools for modeling geometric details. Neural Radiance Field (NeRF) (Mildenhall et al., 2021) represents the 3D scene with implicit continuous functions, mapping point coordinates and view directions to target properties, and employs ray-marching for volume rendering. 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) explicitly initializes a set of Gaussian distributions to represent the geometry and appearance within the scene, improving efficiency and scalability with a rasterization-based rendering pipeline. Omni-Scene (Wei et al., 2024) takes advantage of both pixel and volume-based representations and proposes an Omni-Gaussian representation for improved ego-centric reconstruction and novel view synthesis performance. RenderOcc (Pan et al., 2024) extracts a NeRF-style 3D scene representation and utilizes 2D rendering techniques for 3D supervision from 2D semantics and depth labels. GaussianFormer (Huang et al., 2024b, a) proposes image-to-Gaussian mapping with distribution-based initialization and Gaussian-to-voxel splatting with probabilistic Gaussian superposition for 3D semantic scene completion.
@@ 3.2. Semantic-guided Gaussian Initialization
Design rationale. The large proportion of empty voxels in autonomous driving scenes causes severe spatial redundancy in neural representation techniques like NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023) for modeling non-occupied areas, resulting in high computational expense and slow convergence speed. In light of this, the SGI module first leverages dual-branch 3D representations for global and local semantic information, and then identifies focal voxels as anchors for efficient and effective Gaussian initialization.
Dual-Branch Encoder. Notice that TPV (Huang et al., 2023) representations compress the full voxel representations along three axes respectively, thereby focusing more on global semantics, while dense voxel representations highlight local semantic details. We leverage both voxel and TPV representations for comprehensive aggregation of local and global semantic information. Given the initial 3D features $\textbf{F}_{\rm 3D}$ projected from image features $\textbf{F}_{\rm img}$, the voxel representation is derived through multiple 3D convolution layers followed by a 3D pyramid network:
@@ 3.3. Physical-aware Harmonics Enhancement
Semantic Spherical Harmonics. Spherical Harmonics (SH) demonstrates superior performance in RGB-based color rendering (Kerbl et al., 2023; Wu et al., 2024), showcasing its powerful capability to model spatial variations and complex geometry. To leverage this capability for SSC, we employ the Semantic Spherical Harmonics (SSH), as shown in Figure 4, which incorporates semantic priors into SH basis functions:
@@ 3.4. Training Objective
Following (Cao and De Charette, 2022; Mei et al., 2024), we adopt the cross-entropy loss $\mathcal{L}_{\rm ce}$, lovasz loss $\mathcal{L}_{\rm lovasz}$ (Berman et al., 2018), and scene-class affinity loss $\mathcal{L}_{\rm scal}$ to supervise the final predictions. The overall loss objective is formulated as follows:
@@
Table 1. Camera-based 3D semantic scene completion results on the SemanticKITTI (Behley et al., 2019) validation set. denotes the methods employing EfficientNetB7 (Tan and Le, 2019) as image backbone, and represents the methods utilizing ResNet50 (He et al., 2016) as image backbone. Best results are highlighted in bold, and second-best scores are underlined.
@@
MonoScene (Cao and De Charette, 2022) 36.86 11.08 23.26 0.61 0.45 6.98 1.48 1.86 1.20 0.00 56.52 14.27 26.72 0.46 14.09 5.84 17.89 2.81 29.64 4.14 2.25
TPVFormer (Huang et al., 2023) 35.61 11.36 23.81 0.36 0.05 8.08 4.35 0.51 0.89 0.00 56.50 20.60 25.87 0.85 13.88 5.94 16.92 2.26 30.38 3.14 1.52
OccFormer (Zhang et al., 2023b) 36.50 13.46 25.09 0.81 1.19 25.53 8.52 2.78 2.82 0.00 58.85 19.61 26.88 0.31 14.40 5.61 19.63 3.93 32.62 4.26 2.86
IAMSSC (Xiao et al., 2024) 44.29 12.45 26.26 0.60 0.15 8.74 5.06 1.32 3.46 0.01 54.55 16.02 25.85 0.70 17.38 6.86 24.63 4.95 30.13 6.35 3.56
@@
VoxFormer (Li et al., 2023d) 44.15 13.35 26.54 1.28 0.56 7.26 7.81 1.93 1.97 0.00 53.57 19.69 26.52 0.42 19.54 7.31 26.10 6.10 33.06 9.15 4.94
DepthSSC (Yao and Zhang, 2023) 45.84 13.28 25.94 0.35 1.16 6.02 7.50 2.58 6.32 0.00 55.38 18.76 27.04 0.92 19.23 8.46 26.37 4.52 30.19 7.42 4.09
Symphonies (Jiang et al., 2024) 41.92 14.89 28.68 2.54 2.82 20.44 13.89 3.52 2.24 0.00 56.37 15.28 27.58 0.95 21.64 8.40 25.72 6.60 30.87 9.57 5.76
HASSC (Wang et al., 2024) 44.58 14.74 27.33 1.07 1.14 17.06 8.83 2.25 4.09 0.00 57.23 19.89 29.08 1.26 20.19 7.95 27.01 7.71 33.95 9.20 4.81
H2GFormer (Wang and Tong, 2024) 44.69 14.29 28.21 0.95 0.91 6.80 9.32 1.15 0.10 0.00 57.00 21.74 29.37 0.34 20.51 7.98 27.44 7.80 36.26 9.88 5.81
CGFormer (Yu et al., 2024) 45.99 16.87 34.32 4.61 2.71 19.44 7.67 2.38 4.08 0.00 65.51 20.82 32.31 0.16 23.52 9.20 26.93 8.83 39.54 10.67 7.84
SGN (Mei et al., 2024) 46.21 15.32 33.31 0.61 0.46 6.03 9.84 0.47 0.10 0.00 59.10 19.05 29.41 0.33 25.17 9.96 28.93 9.58 38.12 13.25 7.32
@@ 4.1. Dataset and Evaluation Metrics
SemanticKITTI Benchmark (Behley et al., 2019) is constructed based on the KITTI Odometry Benchmark (Geiger et al., 2012), comprising 22 sequences of real-world autonomous driving scenarios with 20 semantic categories. The semantic occupancy ground truth is generated as $256\times 256\times 32$ voxel grids with a resolution of 0.2mm, within the spatial range of $0\sim 51.2m,-25.6\sim 25.6m,-2\sim 4.4m]$. Following official protocols, sequences 00–07 and 09–10 are designated for training, sequence 08 for validation, and sequences 11–21 for testing.
SSCBench-KITTI-360 Benchmark (Li et al., 2023c) consists of 9 densely annotated sequences in urban driving environments, which are split into a training set with 8,487 frames from sequences (00, 02-05, 07), a validation set with 1,812 frames from sequence (06), and a test set with 2,566 frames from sequence (09). Semantic annotations across 19 classes cover a spatial region of $[0m,51.2m]$ in the forward direction, [25.6m,25.6m][-25.6m,25.6m] laterally, and [2.0m,4.4m][-2.0m,4.4m] vertically, which is uniformly discretized into $256\times 256\times 32$ voxel grids with a resolution of 0.2mm.
Evaluation Metrics. Following (Yu et al., 2024; Mei et al., 2024), we adopt the Intersection over Union (IoU) of occupied voxels as the evaluation metric for the class-agnostic scene completion (SC). We also report the mean Intersection over Union (mIoU) with respect to all semantic classes to measure the performance of the semantic scene completion (SSC).
@@ 4.2. Implementation Details
Network Architecture. Following previous methods (Mei et al., 2024; Li et al., 2023d; Wang et al., 2024), we adopt the ResNet-50 (He et al., 2016) network with FPN (Lin et al., 2017) as the image encoder, generating 2D feature maps with 1/16 of the input resolution. The feature dimension of both 2D and 3D representations $C$ is set to 128. The grid size of 3D feature volume is $(X,Y,Z)=(128,128,16)$, and the final SSC predictions are upsampled to the resolution of $256\times 256\times 32$. In SGI, we employ the dot production to compute voxel-wise feature similarity and select $K=1024$ focal anchors for Gaussian initialization. In the PHE module, we set the regularization parameter for the soft orthogonality loss as $\lambda=1e-6$.
Training Setup. The RGB images from cam2 are cropped into size of $1220\times 370$ as input for the SemanticKITTI dataset, and the RGB images from cam1 are resized to $1408\times 376$ for the SSCBench-KITTI-360 dataset. We train SPHERE for 24 epochs on 4 NVIDIA A6000 GPUs, with a total batch size of 4. About 20 GB of GPU memory is consumed on each GPU while training. We employ a self-distillation training strategy (Wang et al., 2024) for more robust and efficient training process. The AdamW (Loshchilov and Hutter, 2017) optimizer is adopted with an initial learning rate of 2e-4 and a weight decay of 1e-2.
Table 2. Camera-based 3D semantic scene completion results on the SSCBench-KITTI-360 (Behley et al., 2019) test set. denotes the methods employing EfficientNetB7 (Tan and Le, 2019) as image backbone, and represents the methods utilizing ResNet50 (He et al., 2016) as image backbone. Best results are highlighted in bold, and second-best scores are underlined.
@@
MonoScene (Cao and De Charette, 2022) 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
GaussianFormer (Huang et al., 2024b) 35.38 12.92 18.93 1.02 4.62 18.07 7.59 3.35 45.47 10.89 25.03 5.32 28.44 5.68 29.54 8.62 2.99 2.32 9.51 5.14
GaussianFormer2 (Huang et al., 2024a) 38.31 13.90 21.08 2.55 4.21 12.41 5.73 1.59 54.12 11.04 32.31 3.34 32.01 4.98 28.94 17.33 3.57 5.48 5.88 3.54
TPVFormer (Huang et al., 2023) 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer (Zhang et al., 2023b) 40.27 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
IAMSSC (Xiao et al., 2024) 41.80 12.97 18.53 2.45 1.76 5.12 3.92 3.09 47.55 10.56 28.35 4.12 31.53 6.28 29.17 15.24 8.29 7.01 6.35 4.19
@@
VoxFormer (Li et al., 2023d) 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
DepthSSC (Yao and Zhang, 2023) 40.85 14.28 21.90 2.36 4.30 11.51 4.56 2.92 50.88 12.89 30.27 2.49 37.33 5.22 29.61 21.59 5.97 7.77 5.24 3.51
Symphonies (Jiang et al., 2024) 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
CGFormer (Yu et al., 2024) 48.07 20.05 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77
SGN (Mei et al., 2024) 47.06 18.25 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86

### v2

@@ 1. Introduction
In autonomous driving systems, 3D Semantic Scene Completion (SSC)  (roldao20223d) is a crucial task for holistic scene perception, inferring voxel-level geometry and semantics for several downstream tasks. To deal with the inherent complexity of large-scale outdoor scenarios, LiDAR-based SSC methods (roldao2020lmscnet; cheng2021s3cnet; rist2021semantic; yan2021sparse; xia2023scpnet; zhang2023attention; ye2022unsupervised; li2023lode; lee2024semcity) have been the predominant solutions, taking advantage of explicit 3D geometric structural information entailed in the point cloud data. Despite their effectiveness, the inherently high cost and limited scalability of LiDAR sensors constrain practical applications of LiDAR-based methods in real-world deployment. In light of this issue, multi-view images (miao2023occdepth; yao2023ndc; tong2023scene; zheng2024multi) have garnered increasing attention as a more cost-effective and flexible input modality.
The pioneering camera-based SSC method, MonoScene (cao2022monoscene), extracts 2D features from input RGB images and lifts them into dense 3D volume features through depth-based projection. Following this, existing camera-based SSC methods fall mainly into two paradigms: voxel-based and plane-based methods. Voxel-based SSC methods (zhang2023occformer; li2023voxformer) discretize 3D scenes into voxels and assign extracted image features to each voxel, and progress has been achieved in exploiting scene-from-instance feature interactions (jiang2024symphonize), and employing a dense-sparse-dense strategy for hybrid guidance (mei2024camera). In addition, plane-based SSC methods adopt Bird’s-Eye-View (BEV) (huang2021bevdet) and Tri-Perspective View (TPV) (huang2023tri) features for efficient scene representations. Subsequent research further improves the quality of scene representation through camera-aware depth estimation and refinement (li2023bevdepth), dynamic temporal stereo information (li2023bevstereo), and explicit-implicit compact feature projection (ma2024cotr).
While existing SSC methods have demonstrated promising performance in voxel-wise semantic prediction, they frequently fail to capture the underlying physical regularities in the real world, leading to unrealistic geometric details in scene completion results, as illustrated in Figure 1 (a). On the other hand, neural reconstruction methods like NeRF (mildenhall2021nerf) and 3DGS (kerbl20233d) excel at learning realistic geometry details, but suffer from high computational expense and slow convergence in large-scale complex autonomous driving scenes, resulting in confusing semantics, as shown in Figure 1 (b). The above challenges naturally raise a critical question: How can we achieve both accurate semantics and realistic geometry, efficiently and effectively?
@@ 2.1. Camera-based Semantic Scene Completion
Voxel-based SSC methods discretize 3D space into voxels and assign image features to each voxel for scene perception. MonoScene (cao2022monoscene) is a pioneering effort that reconstructs 3D scenes from RGB images by projecting image features to 3D space along optical rays. OccFormer (zhang2023occformer) introduces a dual-path transformer architecture with a mask-wise prediction paradigm, effectively capturing spatial dependencies for improved completion accuracy. VoxFormer (li2023voxformer) presents a two-stage framework that integrates class-agnostic query proposals with class-specific semantic segmentation, facilitating the propagation of discriminative semantic information from seed voxels across the entire scene for enhanced completion. Symphonize (jiang2024symphonize) introduces a scene-from-instance framework that enhances the interplay between image and volumetric representations, enabling more precise scene completion. SGN (mei2024camera) presents a one-stage SSC framework employing a dense-sparse-dense design, which enhances segmentation boundary clarity and improves the accuracy of 3D semantic scene completion. CGFormer (yu2024context) introduces a context- and geometry-aware voxel transformer that utilizes context-dependent queries and multiple 3D representations to enhance scene completion performance.
Plane-based SSC methods compress the 3D space along certain dimensions for efficient view transformation and scene representation. A pioneering approach in this domain is the Lift-Splat-Shoot (LSS) (philion2020lift) framework, which establishes an end-to-end pipeline by first ”lifting” images into feature frustums, then ”splatting” these frustums onto a rasterized BEV grid, and finally ”shooting” template trajectories into a BEV cost map for downstream tasks. Building upon the LSS paradigm, BEVDet (huang2021bevdet) introduces a holistic BEV-based framework for scene understanding, comprising an Image-view Encoder, a View Transformer, a BEV Encoder, and a Task-specific Head. Subsequent achievements have been made in further enhancing the BEV-based representations, incorporating techniques such as inverse perspective mapping (hu2021fiery), angle and radius aware rasterization (liu2023vision), and cross-attention layers with BEV queries (li2022bevformer). BEVDepth (li2023bevdepth) enhances depth learning by introducing a camera-aware depth estimation module alongside a depth refinement module, improving the accuracy of depth predictions. Building on this, BEVStereo (li2023bevstereo) advances depth estimation by incorporating dynamic temporal stereo information. Beyond BEV, TPVFormer (huang2023tri) introduces a Tri-Perspective View (TPV) representation, which decomposes voxel grids into three orthogonal planes, allowing efficient scene encoding.
@@ 2.2. Neural Representation in 3D Perception
Neural representation techniques have emerged as powerful tools for modeling geometric details. Neural Radiance Field (NeRF) (mildenhall2021nerf) represents the 3D scene with implicit continuous functions, mapping point coordinates and view directions to target properties, and employs ray-marching for volume rendering. 3D Gaussian Splatting (3DGS) (kerbl20233d) explicitly initializes a set of Gaussian distributions to represent the geometry and appearance within the scene, improving efficiency and scalability with a rasterization-based rendering pipeline. Omni-Scene (wei2024omni) takes advantage of both pixel and volume-based representations and proposes an Omni-Gaussian representation for improved ego-centric reconstruction and novel view synthesis performance. RenderOcc (pan2024renderocc) extracts a NeRF-style 3D scene representation and utilizes 2D rendering techniques for 3D supervision from 2D semantics and depth labels. GaussianFormer (huang2024gaussianformer; huang2024probabilistic) proposes image-to-Gaussian mapping with distribution-based initialization and Gaussian-to-voxel splatting with probabilistic Gaussian superposition for 3D semantic scene completion.
@@ 3.2. Semantic-guided Gaussian Initialization
Design rationale. The large proportion of empty voxels in autonomous driving scenes causes severe spatial redundancy in neural representation techniques like NeRF (mildenhall2021nerf) and 3DGS (kerbl20233d) for modeling non-occupied areas, resulting in high computational expense and slow convergence speed. In light of this, the SGI module first leverages dual-branch 3D representations for global and local semantic information, and then identifies focal voxels as anchors for efficient and effective Gaussian initialization.
Dual-Branch Encoder. Notice that TPV (huang2023tri) representations compress the full voxel representations along three axes respectively, thereby focusing more on global semantics, while dense voxel representations highlight local semantic details. We leverage both voxel and TPV representations for comprehensive aggregation of local and global semantic information. Given the initial 3D features $\textbf{F}_{\rm 3D}$ projected from image features $\textbf{F}_{\rm img}$, the voxel representation is derived through multiple 3D convolution layers followed by a 3D pyramid network:
@@ 3.3. Physical-aware Harmonics Enhancement
Semantic Spherical Harmonics. Spherical Harmonics (SH) demonstrates superior performance in RGB-based color rendering (kerbl20233d; wu20244d), showcasing its powerful capability to model spatial variations and complex geometry. To leverage this capability for SSC, we employ the Semantic Spherical Harmonics (SSH), as shown in Figure 4, which incorporates semantic priors into SH basis functions:
@@ 3.4. Training Objective
Following (cao2022monoscene; mei2024camera), we adopt the cross-entropy loss $\mathcal{L}_{\rm ce}$, lovasz loss $\mathcal{L}_{\rm lovasz}$ (berman2018lovasz), and scene-class affinity loss $\mathcal{L}_{\rm scal}$ to supervise the final predictions. The overall loss objective is formulated as follows:
@@
Table 1. Camera-based 3D semantic scene completion results on the SemanticKITTI (behley2019semantickitti) validation set. denotes the methods employing EfficientNetB7 (tan2019efficientnet) as image backbone, and represents the methods utilizing ResNet50 (he2016deep) as image backbone. Best results are highlighted in bold, and second-best scores are underlined.
@@
MonoScene (cao2022monoscene) 36.86 11.08 23.26 0.61 0.45 6.98 1.48 1.86 1.20 0.00 56.52 14.27 26.72 0.46 14.09 5.84 17.89 2.81 29.64 4.14 2.25
TPVFormer (huang2023tri) 35.61 11.36 23.81 0.36 0.05 8.08 4.35 0.51 0.89 0.00 56.50 20.60 25.87 0.85 13.88 5.94 16.92 2.26 30.38 3.14 1.52
OccFormer (zhang2023occformer) 36.50 13.46 25.09 0.81 1.19 25.53 8.52 2.78 2.82 0.00 58.85 19.61 26.88 0.31 14.40 5.61 19.63 3.93 32.62 4.26 2.86
IAMSSC (xiao2024instance) 44.29 12.45 26.26 0.60 0.15 8.74 5.06 1.32 3.46 0.01 54.55 16.02 25.85 0.70 17.38 6.86 24.63 4.95 30.13 6.35 3.56
@@
VoxFormer (li2023voxformer) 44.15 13.35 26.54 1.28 0.56 7.26 7.81 1.93 1.97 0.00 53.57 19.69 26.52 0.42 19.54 7.31 26.10 6.10 33.06 9.15 4.94
DepthSSC (yao2023depthssc) 45.84 13.28 25.94 0.35 1.16 6.02 7.50 2.58 6.32 0.00 55.38 18.76 27.04 0.92 19.23 8.46 26.37 4.52 30.19 7.42 4.09
Symphonies (jiang2024symphonize) 41.92 14.89 28.68 2.54 2.82 20.44 13.89 3.52 2.24 0.00 56.37 15.28 27.58 0.95 21.64 8.40 25.72 6.60 30.87 9.57 5.76
HASSC (wang2024not) 44.58 14.74 27.33 1.07 1.14 17.06 8.83 2.25 4.09 0.00 57.23 19.89 29.08 1.26 20.19 7.95 27.01 7.71 33.95 9.20 4.81
H2GFormer (wang2024h2gformer) 44.69 14.29 28.21 0.95 0.91 6.80 9.32 1.15 0.10 0.00 57.00 21.74 29.37 0.34 20.51 7.98 27.44 7.80 36.26 9.88 5.81
CGFormer (yu2024context) 45.99 16.87 34.32 4.61 2.71 19.44 7.67 2.38 4.08 0.00 65.51 20.82 32.31 0.16 23.52 9.20 26.93 8.83 39.54 10.67 7.84
SGN (mei2024camera) 46.21 15.32 33.31 0.61 0.46 6.03 9.84 0.47 0.10 0.00 59.10 19.05 29.41 0.33 25.17 9.96 28.93 9.58 38.12 13.25 7.32
@@ 4.1. Dataset and Evaluation Metrics
SemanticKITTI Benchmark (behley2019semantickitti) is constructed based on the KITTI Odometry Benchmark (geiger2012we), comprising 22 sequences of real-world autonomous driving scenarios with 20 semantic categories. The semantic occupancy ground truth is generated as $256\times 256\times 32$ voxel grids with a resolution of 0.2mm, within the spatial range of $0\sim 51.2m,-25.6\sim 25.6m,-2\sim 4.4m]$. Following official protocols, sequences 00–07 and 09–10 are designated for training, sequence 08 for validation, and sequences 11–21 for testing.
SSCBench-KITTI-360 Benchmark (li2023sscbench) consists of 9 densely annotated sequences in urban driving environments, which are split into a training set with 8,487 frames from sequences (00, 02-05, 07), a validation set with 1,812 frames from sequence (06), and a test set with 2,566 frames from sequence (09). Semantic annotations across 19 classes cover a spatial region of $[0m,51.2m]$ in the forward direction, [25.6m,25.6m][-25.6m,25.6m] laterally, and [2.0m,4.4m][-2.0m,4.4m] vertically, which is uniformly discretized into $256\times 256\times 32$ voxel grids with a resolution of 0.2mm.
Evaluation Metrics. Following (yu2024context; mei2024camera), we adopt the Intersection over Union (IoU) of occupied voxels as the evaluation metric for the class-agnostic scene completion (SC). We also report the mean Intersection over Union (mIoU) with respect to all semantic classes to measure the performance of the semantic scene completion (SSC).
@@ 4.2. Implementation Details
Network Architecture. Following previous methods (mei2024camera; li2023voxformer; wang2024not), we adopt the ResNet-50 (he2016deep) network with FPN (lin2017feature) as the image encoder, generating 2D feature maps with 1/16 of the input resolution. The feature dimension of both 2D and 3D representations $C$ is set to 128. The grid size of 3D feature volume is $(X,Y,Z)=(128,128,16)$, and the final SSC predictions are upsampled to the resolution of $256\times 256\times 32$. In SGI, we employ the dot production to compute voxel-wise feature similarity and select $K=1024$ focal anchors for Gaussian initialization. In the PHE module, we set the regularization parameter for the soft orthogonality loss as $\lambda=1e-6$.
Training Setup. The RGB images from cam2 are cropped into size of $1220\times 370$ as input for the SemanticKITTI dataset, and the RGB images from cam1 are resized to $1408\times 376$ for the SSCBench-KITTI-360 dataset. We train SPHERE for 24 epochs on 4 NVIDIA A6000 GPUs, with a total batch size of 4. About 20 GB of GPU memory is consumed on each GPU while training. We employ a self-distillation training strategy (wang2024not) for more robust and efficient training process. The AdamW (loshchilov2017decoupled) optimizer is adopted with an initial learning rate of 2e-4 and a weight decay of 1e-2.
Table 2. Camera-based 3D semantic scene completion results on the SSCBench-KITTI-360 (behley2019semantickitti) test set. denotes the methods employing EfficientNetB7 (tan2019efficientnet) as image backbone, and represents the methods utilizing ResNet50 (he2016deep) as image backbone. Best results are highlighted in bold, and second-best scores are underlined.
@@
MonoScene (cao2022monoscene) 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
GaussianFormer (huang2024gaussianformer) 35.38 12.92 18.93 1.02 4.62 18.07 7.59 3.35 45.47 10.89 25.03 5.32 28.44 5.68 29.54 8.62 2.99 2.32 9.51 5.14
GaussianFormer2 (huang2024probabilistic) 38.31 13.90 21.08 2.55 4.21 12.41 5.73 1.59 54.12 11.04 32.31 3.34 32.01 4.98 28.94 17.33 3.57 5.48 5.88 3.54
TPVFormer (huang2023tri) 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer (zhang2023occformer) 40.27 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
IAMSSC (xiao2024instance) 41.80 12.97 18.53 2.45 1.76 5.12 3.92 3.09 47.55 10.56 28.35 4.12 31.53 6.28 29.17 15.24 8.29 7.01 6.35 4.19
@@
VoxFormer (li2023voxformer) 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
DepthSSC (yao2023depthssc) 40.85 14.28 21.90 2.36 4.30 11.51 4.56 2.92 50.88 12.89 30.27 2.49 37.33 5.22 29.61 21.59 5.97 7.77 5.24 3.51
Symphonies (jiang2024symphonize) 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
CGFormer (yu2024context) 48.07 20.05 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77
SGN (mei2024camera) 47.06 18.25 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86